OpenStudy (anonymous):

A conductor material has a free electrons density of 10^24 electrons per meter cube. when a voltage s applied a constant drift velocity of 1.5X10^-2 meter/second is attained by the electrons. of the cross sectional area of the material is 1 cm^2, calculate the magnitude of the current. Electronic charge is 1.6X10^-19 coulomb

OpenStudy (anonymous):

Drift velocity is defined by: v = I / nAq v is the drift velocity I is the current n is the density of charge carriers A is the cross-sectional area q is the electric charge magnitude

OpenStudy (anonymous):

after finding the drift velocity, what's next? how can I find the magnitude of the current?

OpenStudy (anonymous):

You can solve this equation for the current: v = I / nAq --> I = v * n * A * q You have all 4 values needed, you just need to multiply them all together to find the current. Be careful with your units! Make sure everything is in meters (i notice the area A is in centimeters^2)

OpenStudy (anonymous):

2400 is the answer I got, is this correct?

OpenStudy (anonymous):

@Pompeii00

OpenStudy (anonymous):

24?

OpenStudy (anonymous):

I = (1.5 x 10^-2) * (10^24) * (0.0001) * (1.6 x 10^-19) I don't have a calculator on me. But if you used the numbers above, then you should have the right answer.

OpenStudy (anonymous):

in converting cm to m. shoukd I consider the square?

OpenStudy (anonymous):

Yes. 1 cm = 0.01 m When squared, you need to conver the square. 1 cm^2 = (0.01)^2 m^2 = 0.0001 m^2

OpenStudy (anonymous):

ok, thank you very much :)